OpenStudy (anonymous):

Can someone help me with this? I need to Find the minimum y-value on a graph of y=f(x) F(x)=x^2+8x-1 What is the minimum y-value? Y=

OpenStudy (tkhunny):

What's the chance you are up on how to "Complete the Square"?

OpenStudy (anonymous):

eh...it confuses the hell outta me

OpenStudy (anonymous):

What class are you looking at this question in? Are you in calculus and familiar with solving extrema problems by taking derivatives?

OpenStudy (tkhunny):

You don't need any of that , anyway. How about -b/(2a)? Ever seen that?

OpenStudy (anonymous):

its algebra yes I have seen that

OpenStudy (tkhunny):

Yikes! I hope you're not in calculus. You better have better algebra before you do that! Completing the Square is important.

OpenStudy (anonymous):

I only have 2 more weeks left of this. I am a human services major, not math - I hate math (and it hates me as well)

OpenStudy (tkhunny):

Oh, perfect. Then we just happen to know that the minimum or maximum value of a parabola is a x = -b/(2a). We have a = 1 and b = 8. Simple as that. Evaluation the function at the resulting value and you'll be done.

OpenStudy (anonymous):

so it is just turned into a different equation replacing the a and b?

OpenStudy (tkhunny):

It's a standard form. \(ax^{2} + bx + c\) Whatever numbers appear, those are their assignments.

OpenStudy (tkhunny):

Where are we? Did you get x = -8/(2*1) = -8/2 = -4? Let's find f(-4) and we'll be done.